# Shell-model study of ${}^{58}$i using quantum computing algorithm

### 23 Feb 2024 v1

@@ Abstract
This study presents a simulated quantum computing approach for the investigation into the shell-model energy levels of ${}^{58}$i through the application of the variational eigensolver (VQE) method in combination with a problem-specific ansatz. The primary objective is to achieve a fully accurate low-lying energy spectrum of ${}^{58}$i. The chosen isotope, ${}^{58}$i is particularly interesting in nuclear physics through its role in astrophysical reactions while also being a simple but not-trivial nucleus for shell-model study, it being two particles outside a closed shell. Our ansatz, along with the VQE method are shown to be able to reproduce exact energy values for the ground state and first and second excited states. We compare a classical shell model code, the values obtained by diagonalization of the Hamiltonian after qubit mapping, and a noiseless simulated ansatz+VQE simulation. The exact agreement between classical and qubit-mapped diagonalisation show the correctness of our method, and the high accuracy of the simulation means that the ansatz is suitable to allow a full reconstruction of the full nuclear wave function.
@@ I Introduction
If quantum computation is to form a future tool for large-scale shell model calculations on near-term hardware, much work will be needed to understand appropriate algorithms to find the eigenstates, and to deal with larger problem sizes. Here we take a step in the latter direction by exploring a heavier system than previously explored, looking at a nucleus (${}^{58}$i) with valence particles in the $fp$-shell. In our exploration, a simulated variational quantum eigensolver (VQE) algorithm is used to compare aspects of the method, in comparison with an exact classical shell model calculation. The choice of ${}^{58}$i is primarily motivated as being the simplest two-valence-particle system in the $fp$ shell, and one with which we can compare our quantum simulation calculations with classical shell model. This isotope is also of particular interest in astrophysics in the $s$-process in AGB stars [27, 28, 29], as a well-studied nuclear for use as a nuclear data benchmark [30], and as a nuclear material [31].
@@ II Theoretical framework
In this study, we give details of our methods to investigate the properties of the ${}^{58}\text{Ni}$ nucleus using quantum computing simulation techniques, along with a classical shell model comparison. We take the model space comprising orbitals $1p_{3/2}$, $0f_{5/2}$, and $1p_{1/2}$ above the ${}^{56}\text{Ni}$ core. This section describes the Hamiltonian employed along with its adaptation for quantum computing simulation. We present the tailored ansatz as appropriate to reach the ground state of the system. Furthermore, we discuss the application of the Variational Quantum Eigensolver (VQE) algorithm and associated optimizers for energy minimization. The implementation of excitation operators, crucial for constructing quantum circuits, is detailed.
@@ II.3 VQE and the wavefunction ansatz
The Variational Quantum Eigensolver (VQE) [36, 37, 38] is a hybrid quantum-classical algorithm proposed as a promising approach for finding ground states in many-body quantum systems. It has been applied in different guises in nuclear physics problems [39, 16, 3, 6, 10, 12] thanks to its ability to be implemented, in some cases, on current quantum computing hardware. The basis of the VQE algorithm is to have a parameterised wave function ansatz on a quantum computer, in which the parameters are free variables; the expectation value of the Hamiltonian in its transformed Pauli matrix form is evaluated on the quantum computer, then a classical algorithm adjusts the parameters of the wave function ansatz to seek a minimum of the Hamiltonian expectation value. This is done through a repeated cycle of expectation value measurements on the quantum computer and directed parameter adjustments on the side of the classical algorithm.
An appropriate ansatz (a guess for the wavefunction) must be chosen. In nuclear physics, this could be inspired by traditional nuclear models or techniques, like the Unitary Coupled Cluster (UCC) method [40, 41], or by problem-specific considerations.
@@
Moving to the first excited state of ${}^{58}$i, with $J=2$, the $M$ value can take values $-2$, $-1$, 00, $1$, or $2$. Opting for simplicity, we choose the $|1,3\rangle\equiv|000000001010\rangle$ combining $|n=1,l=1,j=3/2,m=3/2\rangle$ and $|n=1,l=1,j=3/2,m=1/2\rangle$ as the reference state among all the other possible combinations of qubits for $J$ = 2.
Similarly, for the second excited state with $J=4$, allowing $M$ to take values like $-4$, $-3$, $-2$, $-1$, 00, $1$, $2$, $3$, or $4$, we choose the extremal value to avoid contamination with lower $J$ contributions. Here, we have only two combination for $M=4$, and we opt for the $|1,5\rangle\equiv|00000100010\rangle$ as the reference state, corresponding to states $|n=1,l=1,j=3/2,m=3/2\rangle$ and $|n=0,l=3,j=5/2,m=5/2\rangle$.
Having created a simple single configuration reference state for a given target state, parameterised operators are applied to explore the other configurations expected to contribute to the given M-valued wave function. We use parameterized Givens rotation gates [42] to access selected configurations in ${}^{58}$i with correct $M$ values. These gates are particle number preserving, keeping us in the correct Fock space, and are suitably parameterised with enough freedom to find the lowest energy configuration. The excitations in the quantum simulations have been made using the single and double Givens excitation gates. Single excitations, shown in Figure 1, involve the application of creation and annihilation operators to excite a single particle from one orbital to another, as represented by $a^{\dagger}_{i}a_{j}$. Since the nucleon carries an intrinsic angular momentum (spin), the single excitation process is applied so that it conserves $m_{\alpha}$ by choosing the individual spins of the nucleons involved in the excitation.
@@
The prepared ansatz for the ground state, first excited state, and second excited state is represented in Figure 3. Note that while all qubits are active in the ground state wave function, only three are active in the $J=4$ state as only they can give a total projection $M=4$, while nine are active for the $J=2$ state.

### 29 Feb 2024 v2

@@ Abstract
This study presents a simulated quantum computing approach for the investigation into the shell-model energy levels of ${}^{58}$i through the application of the variational eigensolver (VQE) method in combination with a problem-specific ansatz. The primary objective is to achieve a fully accurate low-lying energy spectrum of ${}^{58}$i. The chosen isotope, ${}^{58}$i is particularly interesting in nuclear physics through its role in astrophysical reactions while also being a simple but not-trivial nucleus for shell-model study, it being two particles outside a closed shell. Our ansatz, along with the VQE method are shown to be able to reproduce exact energy values for the ground state and first and second excited states. We compare a classical shell model code, the values obtained by diagonalization of the Hamiltonian after qubit mapping, and a noiseless simulated ansatz+VQE simulation. The exact agreement between classical and qubit-mapped diagonalization shows the correctness of our method, and the high accuracy of the simulation means that the ansatz is suitable to allow a full reconstruction of the full nuclear wave function.
@@ I Introduction
If quantum computation is to form a future tool for large-scale shell model calculations on near-term hardware, much work will be needed to understand appropriate algorithms to find the eigenstates and to deal with larger problem sizes. Here we take a step in the latter direction by exploring a heavier system than previously explored, looking at a nucleus (${}^{58}$i) with valence particles in the $fp$-shell. In our exploration, a simulated variational quantum eigensolver (VQE) algorithm is used to compare aspects of the method, in comparison with an exact classical shell model calculation. The choice of ${}^{58}$i is primarily motivated as being the simplest two-valence-particle system in the $fp$ shell, and one with which we can compare our quantum simulation calculations with the classical shell model. This isotope is also of particular interest in astrophysics in the $s$-process in AGB stars [27, 28, 29], as a well-studied nuclear for use as a nuclear data benchmark [30], and as a nuclear material [31].
@@ II Theoretical framework
In this study, we give details of our methods to investigate the properties of the ${}^{58}$i nucleus using quantum computing simulation techniques, along with a classical shell model comparison. We take the model space comprising orbitals $1p_{3/2}$, $0f_{5/2}$, and $1p_{1/2}$ above the ${}^{56}$i core. This section describes the Hamiltonian employed along with its adaptation for quantum computing simulation. We present the tailored ansatz as appropriate to reach the ground state of the system. Furthermore, we discuss the application of the Variational Quantum Eigensolver (VQE) algorithm and associated optimizers for energy minimization. The implementation of excitation operators, crucial for constructing quantum circuits, is detailed.
@@ II.3 VQE and the wavefunction ansatz
The Variational Quantum Eigensolver (VQE) [36, 37, 38] is a hybrid quantum-classical algorithm proposed as a promising approach for finding ground states in many-body quantum systems. It has been applied in different guises in nuclear physics problems [39, 16, 3, 6, 10, 12] thanks to its ability to be implemented, in some cases, on current quantum computing hardware. The basis of the VQE algorithm is to have a parameterized wave function ansatz on a quantum computer, in which the parameters are free variables; the expectation value of the Hamiltonian in its transformed Pauli matrix form is evaluated on the quantum computer, then a classical algorithm adjusts the parameters of the wave function ansatz to seek a minimum of the Hamiltonian expectation value. This is done through a repeated cycle of expectation value measurements on the quantum computer and directed parameter adjustments on the side of the classical algorithm. An appropriate ansatz (a guess for the wave function) must be chosen. In nuclear physics, this could be inspired by traditional nuclear models or techniques, like the Unitary Coupled Cluster (UCC) method [40, 41], or by problem-specific considerations.
@@
Moving to the first excited state of ${}^{58}$i, with $J=2$, the $M$ value can take values $-2$, $-1$, 00, $1$, or $2$. Opting for simplicity, we choose the $|1,3\rangle\equiv|000000001010\rangle$ combining $|n=1,l=1,j=3/2,m=3/2\rangle$ and $|n=1,l=1,j=3/2,m=1/2\rangle$ as the reference state among all the other possible combinations of qubits for $J$ = 2. Similarly, for the second excited state with $J=4$, allowing $M$ to take values like $-4$, $-3$, $-2$, $-1$, 00, $1$, $2$, $3$, or $4$, we choose the extremal value to avoid contamination with lower $J$ contributions. Here, we have only two combination for $M=4$, and we opt for the $|1,5\rangle\equiv|00000100010\rangle$ as the reference state, corresponding to states $|n=1,l=1,j=3/2,m=3/2\rangle$ and $|n=0,l=3,j=5/2,m=5/2\rangle$.
Having created a simple single configuration reference state for a given target state, parameterized operators are applied to explore the other configurations expected to contribute to the given $M$-valued wave function. We use parameterized Givens rotation gates [42] to access selected configurations in ${}^{58}$i with correct $M$ values. These gates are particle number preserving, keeping us in the correct Fock space, and are suitably parameterized with enough freedom to find the lowest energy configuration. The excitations in the quantum simulations have been made using the single and double Givens excitation gates.
Single excitations, shown in Figure 1, involve the application of creation and annihilation operators to excite a single particle from one orbital to another, as represented by $a^{\dagger}_{i}a_{j}$. Since the nucleon carries an intrinsic angular momentum (spin), the single excitation process is applied so that it conserves $m_{\alpha}$ by choosing the individual spins of the nucleons involved in the excitation.
@@
The prepared ansatz for the ground state, first excited state, and second excited state is represented in Figure 3. Note that while all qubits are active in the ground state wave function, only three are active in the $J=4$ state as only they can give a total projection $M=4$, while nine are active in the $J=2$ state.